Question 16, chap 132, sect 3. part 1 of 2 10 points In an RL circuit, the electromotive force is 5 V, the inductance is 32.7 mH, and the resistance is 7 ?.
Answer #1

part 1:

assuming the inductor to be fully discharged before the switch is closed, equation of current as a function of time is given as:

i(t)=I0*(1-exp(-t/time constant))

where I0=saturated current=emf/resistance

=5/7=0.7143 A

time constant=L/R

=0.0046714 seconds

so at 0.7 ms, current=0.7143*(1-exp(-0.7*0.001/0.0046714))

=0.1 A

part 2:

maximum value of current will happen when inductor is saturated and behave as short circuit,

at that point, resistance in the circuit=7 ohms

so current=5/7=0.7143 A